JoelTheBoss (joel_the_boss):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of two thirds? 2x + 3y = 3 2x − 3y = 3 3x + 2y = 3 3x − 2y = 3 @DolphinDawg @thatgirlcaleigh

OpenStudy (anonymous):

The point at which the line passes through is (6,3) i.e. x1 = 6 & y1 = 3, and it has a slope or gradient, m = 2/3. In order to get the line that passes through those points, you need to make use of the formula: y - y1 = m(x - x1), by substitution you'll have y - 3 = (2/3)(x - 6) 3(y - 3) = 2(x - 6) 3y - 9 = 2x - 12 3y - 2x = - 12 + 9 3y - 2x = - 3 2x - 3y = 3 ...Ans. Option B is correct.